Find two negative consecutive odd integers such that their product is 63.
Answer by edjones(8007)

Let the integers be x and x+2
x(x+2)=63
x^2+2x=63
x^2+2x-63=0
(x+9)(x-7)=0
x=-9
-9+2=-7 (tricky)
Check:
-9*-7=63
Ed